[Brahms-l] Shift report 20040326 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Fri Mar 26 2004 - 15:59:45 EST
------------------------------------------------------------
Date: Fri Mar 26 15:59:45 2004, Shift: 08:00-16:00
Supervisor: sjs, with: 
Log book#15 pages 99 - 99

Shift Summary:
------------------------------------------------------------
1. After early problems, one good store during shift.
2. 
   Pedestal runs 11317, 11320
   3 deg 1/2 field; 40 deg 1050:
      runs 11323, 11325, 11326 

   Dana has fixed the wiring on TOFW 110
   Shift instructions have been updated (pdf file as well...)
3. No new problems found.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Mar 26 16:00:14 2004

This archive was generated by hypermail 2.1.8 : Fri Mar 26 2004 - 16:00:27 EST